If the sum of the measure if the interior angles of a polygon equals the sum of the measure of its exterior angles how many sides does it have?
Answer by jim_thompson5910(35256)

180(n-2) = 360


n-2 = 360/180


n-2 = 2


n = 2+2


n = 4


So there are 4 sides